OpenStudy (rissyroo118):

Which is the equation in slope-intercept form for the line that passes through (2, 8) and is perpendicular to 4x + 2y = 6?

OpenStudy (imstuck):

Ok, the rule about lines being perpendicular to one another is that their slopes are opposite reciprocals of one another. This means that if one line has a slope of 2, the line that runs perpendicular to it will have a slope of -1/2. Does that sound familiar?

OpenStudy (jdoe0001):

what's the slope of -> 4x + 2y = 6?

OpenStudy (imstuck):

Ok, first let's find the slope of the equation. Remember you will solve for y here. Get y alone. 2y = -4x + 6. Solve for y: y = -4/2x + 6. The -4/2 will reduce to -2. So the slope of that line is -2. The opposite reciprocal of - 2 is a positive 1/2 (the reciprocal of 2 is 1 over 2). Now that we have a slope value (which is denoted as "m", don't ask me why), we can use the point we have in a point-slope formula to find the equation.

OpenStudy (rissyroo118):

a y=-1/2x+7 b y=1/2x+7 c y= -2x+12 d y=2x+4

OpenStudy (imstuck):

The point-slope formula looks like this\[y - y _{1}=m(x-x _{1})\]where the y1 and the x1 are coordinates from your given point of (2, 8). 2 is the x coordinate and 8 is the y coordinate. \[y - 8 = \frac{ 1 }{ 2 }(x-2)\]Now wed distribute the 1/2 into the paranthesis to get\[y-8=\frac{ 1 }{ 2 }x-\frac{ 2 }{ 2 }\]which is actually \[y-8=\frac{ 1 }{ 2 }x-1\]Now add the 8 that is with the y to both sides to get \[y=\frac{ 1 }{ 2 }x+7\]That's your equation that is perpendicular to the one they gave you

OpenStudy (rissyroo118):

now - Which is the equation of a line perpendicular to the line with the equation 3x + 4y = 7? 4x + 3y = 3 4x – 3y = 5 3x – 4y = 3 3x + 4y = 5

OpenStudy (rissyroo118):

@IMStuck

OpenStudy (imstuck):

I'm back now!

OpenStudy (rissyroo118):

i figured that one out. its b (:

OpenStudy (imstuck):

First solve it for y. Move the 3x over to the other side using subtraction. 4y = -3x + 7. Now you have to divide everything by 4 to get the y all alone (y's appear to be very lonely, don't they?). When you do that you get y = -3/4x + 7/4. The slope here is -3/4. The opposite reciprocal is 4/3 (positive). Now let's test our choices for the one that gives us a slope of 4/3.

OpenStudy (imstuck):

Oh, you're right! It is B!!! Good!